a camp had enough food to last 50 days for 2200 soldiers . after 17 days some more soldiers joined the camp and the food remaining lasts for 20 days find the number of soldiers who joined later

Answer by ankor@dixie-net.com(22740)

a camp had enough food to last 50 days for 2200 soldiers.
50 * 2200 = 110000 man-food days available
:
After 17 days some more soldiers joined the camp and the food remaining lasts for 20 days
Find the number of soldiers who joined later
Let x = no. of soldiers who joined later
:
17(2200) + 20(x+2200) = 110000
37400 + 20x + 44000 = 110000
20x + 81400 = 110000
20x = 110000 - 81400
20x = 28600
x = 28600/20
x = 1430 men joined after 17 days
: